Almerinda Company estimates that total factory overhead costs will be $2,124,000 for the year. Direct labor hours are estimated to be 531,000.

Required:

A. For Almerinda Company, determine the predetermined factory overhead rate using direct labor hours as the activity base.
B. During April, Almerinda Company accumulated 18,400 hours of direct labor costs on Job 50 and 23,200 hours on Job 51. Determine the amount of factory overhead applied to Jobs 50 and 51 in April.
C. Prepare the journal entry on Apr. 30 to apply factory overhead to both jobs in April according to the predetermined overhead rate. Refer to the Chart of Accounts for exact wording of account titles.
Answer #1

a) Overhead rate = estimated overhead/estimated labor hour = 2124000/531000 = 4 per labor hour

b) Applied overhead

Job 50 = 18400*4 = 73600

Job 51 = 23200*4 = 92800

c) Journal entry

Date account and explanation Debit Credit
Work in process (73600+92800) 166400
Manufacturing overhead 166400
(To record applied overhead)